# A Broad Comparative Evaluation of Software Debloating Tools

### v2

@@ Abstract
Software debloating tools seek to improve the program security and performance by removing unnecessary code, called bloat. While many techniques have been proposed, several barriers to their adoption have emerged. Namely, debloating tools are highly specialized, making it difficult for adopters to find the right type of tool for their needs. This is further hindered by a lack of established metrics and comparative evaluations between tools. To close this information gap, we surveyed 10 years of debloating literature and several tools currently under commercial development to taxonomize knowledge about the debloating ecosystem. We then conducted a broad comparative evaluation of 10 debloating tools to determine their relative strengths and weaknesses. Our evaluation, conducted on a diverse set of 20 benchmark programs, measures tools across 12 performance, security, and correctness metrics.
Our evaluation surfaces several concerning findings that contradict the prevailing narrative in debloating literature. First, debloating tools lack the required maturity to be used on real-world software, evidenced by a slim 21% overall success rate for creating passable debloated versions of medium- and high-complexity benchmarks. Second, debloating tools struggle to produce sound and robust programs. Using our novel differential fuzzing tool, DIFFER, we discovered that only 13% of our debloating attempts produced a sound and robust debloated program. Finally, our results indicate that debloating tools typically do not improve the performance or security posture of debloated programs by a significant degree. We believe that our contributions in this paper will help potential adopters better understand the landscape of tools and will motivate future research and development of more capable debloating tools. To this end, we have made our benchmark set, data, and custom tools publicly available.
@@ 1 Introduction
First, debloating tools are highly specialized. Each tool targets a different kind of bloat, type of software (compiled, interpreted, kernels, etc.), stage of the life cycle (source, IR, binary, etc.) and has a unique analysis method. This makes it challenging for adopters to select the right category of tools for their specific problem. Second, the literature lacks established metrics, resulting in unclear, incomplete, and potentially misleading claims regarding how different methods improve performance and security. This occurs because authors borrow or create their own metrics, which has been shown to be problematic [12, 51]. Ultimately, it is difficult for adopters to understand what benefits to expect as it is not immediately apparent which metrics are useful or relevant. This is exacerbated by superficial explorations of soundness risks in the literature, making it similarly difficult for adopters to understand debloating’s trade-offs and consequences. Finally, the few comparative evaluations present in the literature are limited to a small number of tools, making it difficult for adopters to determine which tools are best in class.
While broad studies on bloat prevalence [53, 54, 14], debloating trade-offs [76], and security metrics [12, 8, 21] have been published, the largest comparative evaluation of software debloating tools to date [4] is limited to only four tools, each of a separate type. This permits comparing individual tools across categories and how well they compose, however the quantity and diversity of tools is not sufficient to compare tools within categories or draw conclusions about the relative merits of different types of debloaters to each other. Second, the evaluation uses only four metrics and may fail to detect important negative impacts of debloating such as excessive runtime or memory consumption in debloated programs.
@@ 2.2 Debloating Workflow and Use Cases
    Specification: The user creates a tool-specific specification that outlines what program behaviors should be kept or eliminated. [^1] Some tools targeting Type I bloat only may not require specifications.
@@ 2.3.2 Compiler-Based Specializers (CBS)
Because CBS are limited to aggressive use cases, they generally only require the user to know how the program will be executed on their system (i.e., a single command line invocation). As a result, creating debloating specifications for CBS typically takes less than an hour. However, there are some notable exceptions that may increase the amount of experience or time needed to configure and use the tool. Since the tools are tightly coupled with the compiler, in some cases the user may need to have working knowledge of the tools build configuration (i.e., compiler flags). Additionally, some tools may require the user to perform manual steps in some cases, such as identifying the boundary between input parsing and program logic if it cannot be automatically identified[3].
@@ 2.3.3 Binary-to-Binary (B2B) Debloaters
B2B debloaters generally require working technical knowledge of the target program only. Like test case-based S2S debloaters, creating a specification requires a large number of test cases in practice to minimize soundness issues. For more complex programs, this may take users one or more hours to enumerate for non-trivial programs. Once processed, debloated binaries are difficult to maintain or further modify due to the limits of the transformation techniques mentioned previously. However, given that B2B debloaters are primarily used for legacy programs, this may not be a significant detriment to their use.
@@ 3.1 Tool Selection
Due to the size and diversity of the software debloating ecosystem, we first scoped our evaluation to tools that support userland C/C++ programs and libraries for x86/x86-64 machines. This benchmark profile is the most widely supported, and had the largest pool of candidate tools. In total, we identified 31 candidate tools after filtering successor tools (e.g., OCCAMv2[45] versus OCCAMv1 [42]). We successfully obtained source code via a public repository or request to the authors for 24 tools, and were able to successfully build and run 17 of them. We expended significant effort to resolve issues for the seven failed tools, however we were not successful due to irreconcilable technical issues [55, 9, 75, 47] or unresponsive authors[64, 77, 15]. Of the remaining 17 tools, we further excluded one because it requires a commercial license for IDA Pro [54] and six others because they were not readily adaptable to new targets. The underlying causes for exclusion included one tool that was hardcoded to its benchmarks[72], several tools that required time consuming and manual pre-processing steps for new benchmarks [11, 50, 49, 76], and one tool with technical limitations [22].
@@ 4.1 Tool Performance
With our testing environment configured and manual harnessing complete, we then ran each tool ten times [^3] Due to prohibitively long run times on the order of several CPU hours/days (even with our customization), we ran several tools on certain benchmarks only once: 1) all CHISEL and CHISEL-GT performance tests, 2) OCCAM and TRIMMER on objdump. on each compatible benchmark program recording the total CPU runtime and peak memory usage per run. We summarize the average run time in CPU minutes and peak memory use across all benchmarks at each level of complexity in Table 4. Note that we include debloating operations that fail in these calculations. We exclude only benchmarks that are not compatible with the tool for reasons such as the tool does not support C++ code, multi-threaded programs, etc. The first set of columns in Table 4 indicate the number of compatible benchmarks per tool at each level of complexity, out of the total 10 low, 6 medium, and 4 high complexity benchmarks.
@@ 4.2 Program Performance
Our evaluation reveals several interesting results. RAZOR and Libfilter produced debloated binaries that are larger on average across all complexity levels. This is not surprising, as RAZOR stitches the debloated version of the program into a new code section of the original binary and marks the original code section as non-executable rather than actually removing it. Similarly, Libfilter rewrites unreachable functions as HLT instructions rather than removing them. We note that reducing binary size is not a stated goal of either tool, and this metric was not evaluated by their authors. Further, the transformation methods used by these tools can be replaced with other methods that reduce binary size rather than increase it without fundamental changes to their approaches.
@@
With respect to our performance metrics, we observed only a few instances where debloated programs performed noticeably worse than their unmodified counterparts. Excluding an outlier for CHISEL, only TRIMMER and OCCAM demonstrated consistent and significant (>4%) negative impacts to run time or memory consumption across multiple complexity levels. The most significant of which are large (28% to 76%) increases in run time. This is somewhat unsurprising as these debloaters significantly alter the build chain and overall phase ordering of compiler optimizations to facilitate debloating. Phase ordering has significant effects on the binary performance and is an open area of research[38, 39].
@@ 4.3 Program Security Posture
Special Purpose Gadget Availability. The third code re-usability metrics we evaluated is concerned with the availability of special purpose code reuse gadgets in a program binary. Attackers use special purpose gadgets as scaffolding to organize functional gadgets or to accomplish specific tasks such as execute system calls (i.e., syscall gadget). The special purpose gadget availability metric captures how many different types of special purpose gadget types are available to an attacker, defined as the presence of at least one gadget per type. GSA defines 10 different types of special purpose gadgets, the most important of these types being syscall gadgets. Syscall gadgets are used to invoke operating system functions such as exec or mprotect on the attacker’s behalf. These gadgets are rare outside of libc, as they only appear unintentionally in typical userland programs.
@@
Overall, debloating did not have a significant impact on the types of special purpose gadgets available; on average the total number of types was increased or decreased by less than one type for almost all tools and complexity levels. CHISEL-GT and TRIMMER both introduced more than one type of special purpose gadget on average for medium+high and high complexity benchmarks, respectively. LMCAS-SIFT and Binary Reduce (Static) both reduce the availability of special purpose gadget types significantly at several levels of benchmark complexity, although it is important to note that this is due to the elimination of large amounts of code in libraries via static linking. Focusing specifically on syscall gadgets, the results were similarly mixed. Most tools either eliminated syscall gadgets for some benchmarks, introduced them in others, or more commonly, both.
@@ 4.4 Program Correctness / Robustness
Debloating transformations are complex, and ensuring they do not compromise the soundness of the program is challenging whether they are implemented pre- (i.e., S2S), during (i.e., CBS), or post-production (i.e., B2B, SL). Flawed or incomplete transformations have the potential to negatively impact programs in many ways that can manifest as logic bugs, run-time errors, crashes, and in some cases introducing new vulnerabilities. We discovered ample evidence of this during our benchmark performance tests (Section 4.2), where 36.6% of the debloated binaries produced in our evaluation failed to execute their retained functionality.
@@
It is important to note that DIFFER does not and cannot provide formal guarantees of soundness in debloaters or the debloated programs they produce. Like other dynamic analysis testing approaches, DIFFER cannot exhaustively test the input space for complex programs in the general case. Still, DIFFER is quite useful for post-debloating validation as it is user friendly and only requires a moderate about of user expertise to configure.
Using the general debloating specifications we created in Section 3.3 as a starting point, we configured DIFFER to test both retained and debloated functions for each benchmark program. We ran DIFFER on each of the debloated benchmark programs that successfully completed their performance test (a total of 90 variants) for a maximum of 12 hours to identify crashes, inconsistencies, or errors introduced during debloating as well as failures to remove functionality marked for debloating. Our results are concerning: after filtering for false positives, DIFFER discovered errors or crashes in 27.8% (25 of 90) of debloated benchmarks. Additionally, debloating tools failed to remove features marked for debloating in 60% (39 of 65) of the remaining debloated benchmarks. Of the final 26 debloated benchmarks that passed testing by DIFFER, thirteen were produced by Binary Reduce (static), seven were produced by libfilter, four were produced by LMCAS-SIFT, and OCCAM and CHISEL each produced one. Across our evaluation, 50% (20 of 40) of attempts to remove Type I bloat were ultimately successful, and only 3.3% of attempts to remove Type II bloat were successful (6 of 180) [^4] Five successes were on low-complexity benchmarks, one was on a medium complexity benchmark.. Our results indicate that the general lack of attention paid to post-debloating validation has resulted in over-reporting of successful debloating in the literature.
@@ 5 Discussion and Key Findings
Tool Maturity. Our most important finding is that debloating tools lack the maturity required to be used on real-world software. This is evidenced by a slim success rate of 42.5% across all tools and complexity levels at producing debloated programs or libraries that can pass a performance test. The success rate across medium- and high-complexity benchmarks is only 21%, indicating that the current generation of debloaters are suitable for use on lower complexity programs only. To advance the state of the art in debloating, new approaches should focus on supporting more complex software packages and programming language features / paradigms.
Soundness Issues. When tested for correctness and robustness with DIFFER, only two of the ten tools we evaluated, Binary Reduce static and dynamic, did not produce unsound debloated binaries (i.e., DIFFER did not detect any crashes, failures, or errors in kept functionality). In the case of Binary Reduce (dynamic), this tool produced only one debloated benchmark, and DIFFER did detect that it still contained features marked for debloating. As such, it is possible this binary is only sound because the tool did not successfully debloat any features, which is supported by the fact that this binary increased in size after debloating.
@@
Marginal Performance and Security Benefits. Finally, the tools we evaluated demonstrated consistent and significant improvements in two metrics only: binary size and gadget locality. We observed marginal benefits at best for all other performance and security metrics. It is important to note that our security evaluation is limited to code re-usability metrics due to the difficulty of quantifying security improvements. These tools may have had positive impacts such as removing previously-unknown reachable vulnerabilities that are generally not possible to evaluate because they are undecidable.
Considering these marginal benefits in concert with the aforementioned maturity and soundness issues, users are likely to find the benefits of software debloating are currently outweighed by the costs and potential risks in real-world scenarios.

### v3

@@ Abstract
Software debloating tools seek to improve program security and performance by removing unnecessary code, called bloat. While many techniques have been proposed, several barriers to their adoption have emerged. Namely, debloating tools are highly specialized, making it difficult for adopters to find the right type of tool for their needs. This is further hindered by a lack of established metrics and comparative evaluations between tools. To close this information gap, we surveyed 10 years of debloating literature and several tools currently under commercial development to taxonomize knowledge about the debloating ecosystem. We then conducted a broad comparative evaluation of 10 debloating tools to determine their relative strengths and weaknesses. Our evaluation, conducted on a diverse set of 20 benchmark programs, measures tools across 12 performance, security, and correctness metrics.
Our evaluation surfaces several concerning findings that contradict the prevailing narrative in the debloating literature. First, debloating tools lack the maturity required to be used on real-world software, evidenced by a slim 22% overall success rate for creating passable debloated versions of medium- and high-complexity benchmarks. Second, debloating tools struggle to produce sound and robust programs. Using our novel differential fuzzing tool, DIFFER, we discovered that only 13% of our debloating attempts produced a sound and robust debloated program. Finally, our results indicate that debloating tools typically do not improve the performance or security posture of debloated programs by a significant degree according to our evaluation metrics. We believe that our contributions in this paper will help potential adopters better understand the landscape of tools and will motivate future research and development of more capable debloating tools. To this end, we have made our benchmark set, data, and custom tools publicly available.
@@ 1 Introduction
First, debloating tools are highly specialized. Each tool targets a different kind of bloat, type of software (compiled, interpreted, kernels, etc.), stage of the life cycle (source, IR, binary, etc.) and has a unique analysis method. This makes it challenging for adopters to select the right category of tools for their specific problem. Second, the literature lacks established metrics resulting in unclear, incomplete, and potentially misleading claims regarding how different methods improve performance and security. This occurs because authors borrow or create their own metrics, which has been shown to be problematic [12, 51]. Ultimately, it is difficult for adopters to understand what benefits to expect as it is not immediately apparent which metrics are useful or relevant. This is exacerbated by superficial explorations of soundness risks in the literature, making it similarly difficult for adopters to understand debloating’s trade-offs and consequences. Finally, the few comparative evaluations present in the literature are limited to a small number of tools, making it difficult for adopters to determine which tools are best in class.
While broad studies on bloat prevalence [53, 54, 14], debloating trade-offs [76], and security metrics [12, 8, 21] have been published, the largest comparative evaluation of software debloating tools to date [4] is limited to only four tools, each of a separate type. This permits comparing individual tools across categories and how well they compose, however the quantity and diversity of tools is not sufficient to compare tools within categories or draw conclusions about the relative merits of different types of debloaters. Also, the evaluation uses only four metrics and may fail to detect important negative impacts of debloating such as excessive runtime or memory consumption in debloated programs.
@@ 2.2 Debloating Workflow and Use Cases
    Specification: The user creates a tool-specific specification that outlines what program behaviors should be kept or eliminated. [^1] Tools targeting only Type I bloat may not require specifications.
@@ 2.3.2 Compiler-Based Specializers (CBS)
Because CBS are limited to aggressive use cases, they generally only require the user to know how the program will be executed on their system (i.e., a single command line invocation). As a result, creating debloating specifications for CBS typically takes less than an hour. However, there are some notable exceptions that may increase the amount of experience or time needed to configure and use the tool. Since the tools are tightly coupled with the compiler, in some cases the user may need to have working knowledge of the tools build configuration (i.e., compiler flags). Additionally, some tools may require the user to perform manual steps, such as identifying the boundary between input parsing and program logic if it cannot be automatically identified[3].
@@ 2.3.3 Binary-to-Binary (B2B) Debloaters
B2B debloaters generally require working technical knowledge of the target program only. Like test case-based S2S debloaters, creating a specification requires a large number of test cases in practice to minimize soundness issues. For more complex programs, this may take users one or more hours to enumerate. Once processed, debloated binaries are difficult to maintain or further modify due to the limits of the transformation techniques mentioned previously. However, given that B2B debloaters are primarily used for legacy programs, this may not be a significant detriment to their use.
@@ 3.1 Tool Selection
Due to the size and diversity of the software debloating ecosystem, we first scoped our evaluation to tools that support userland C/C++ programs and libraries for x86/x86-64 machines. This benchmark profile is the most widely supported and had the largest pool of candidate tools. In total, we identified 31 candidate tools after filtering predecessor tools (e.g., including OCCAMv2[45] versus OCCAMv1 [42]). We obtained source code via a public repository or request to the authors for 24 tools, and were able to successfully build and run 17 of them. We expended significant effort to resolve issues for the seven failing builds, however we were not successful due to irreconcilable technical issues [55, 9, 75, 47] or unresponsive authors[64, 77, 15]. Of the remaining 17 tools, we further excluded one because it requires a commercial license for IDA Pro [54] and six others because they were not readily adaptable to new targets. The underlying causes for exclusion included one tool that was hardcoded to its benchmarks[72], several tools that required time consuming and manual pre-processing steps for new benchmarks [11, 50, 49, 76], and one tool with technical limitations [22].
@@ 4.1 Tool Performance
With our testing environment configured and manual harnessing complete, we then ran each tool ten times [^3] Due to prohibitively long run times on the order of several CPU hours/days (even with customization), we ran several tools on certain benchmarks only once: 1) all CHISEL and CHISEL-GT performance tests, 2) OCCAM and TRIMMER on objdump. on each compatible benchmark program recording the total CPU runtime and peak memory usage per run. We summarize the average run time in CPU minutes and peak memory use across all benchmarks at each level of complexity in Table 4. Note that we include debloating operations that fail in these calculations. We exclude only benchmarks that are not compatible with the tool for reasons such as the tool does not support C++ code, multi-threaded programs, etc. The first set of columns in Table 4 indicate the number of compatible benchmarks per tool at each level of complexity, out of the total 10 low, 6 medium, and 4 high complexity benchmarks.
@@ 4.2 Program Performance
Our evaluation reveals several interesting results. RAZOR and Libfilter produced debloated binaries that are larger on average across all complexity levels. This is not surprising, as RAZOR stitches the debloated version of the program into a new code section of the original binary and marks the original code section as non-executable rather than actually removing it. Similarly, Libfilter rewrites unreachable functions as HLT (halt) instructions rather than removing them. We note that reducing binary size is not a stated goal of either tool, and this metric was not evaluated by their authors. Further, the transformation methods used by these tools can be replaced with other methods that reduce binary size rather than increase it without fundamental changes to their approaches.
@@
With respect to our performance metrics, we observed only a few instances where debloated programs performed noticeably worse than their unmodified counterparts. Excluding an outlier for CHISEL, only TRIMMER and OCCAM demonstrated consistent and significant (>4%) negative impacts to run time or memory consumption across multiple complexity levels. The most significant of which are large (28% to 76%) increases in run time. This is somewhat unsurprising as these debloaters significantly alter the build chain and overall phase ordering of compiler optimizations to facilitate debloating. Phase ordering has significant effects on binary performance and is an open area of research[38, 39].
@@ 4.3 Program Security Posture
Special Purpose Gadget Availability. The third code re-usability metric we evaluated is concerned with the availability of special purpose code reuse gadgets in a program binary. Attackers use special purpose gadgets as scaffolding to organize functional gadgets or to accomplish specific tasks such as execute system calls (i.e., syscall gadget). The special purpose gadget availability metric captures how many different types of special purpose gadget types are available to an attacker, defined as the presence of at least one gadget per type. GSA defines 10 different types of special purpose gadgets, the most important of these types being syscall gadgets. Syscall gadgets are used to invoke operating system functions such as exec or mprotect on the attacker’s behalf. These gadgets are rare outside of libc, as they only appear unintentionally in typical userland programs.
@@
Overall, debloating did not have a significant impact on the types of special purpose gadgets available. On average, the total number of types was increased or decreased by less than one type for almost all tools and complexity levels. CHISEL-GT and TRIMMER both introduced more than one type of special purpose gadget on average for medium+high and high complexity benchmarks, respectively. LMCAS-SIFT and Binary Reduce (Static) both reduce the availability of special purpose gadget types significantly at several levels of benchmark complexity, although it is important to note that this is due to the elimination of large amounts of code in libraries via static linking. Focusing specifically on syscall gadgets, the results were similarly mixed. Most tools either eliminated syscall gadgets for some benchmarks, introduced them in others, or more commonly, both.
@@ 4.4 Program Correctness / Robustness
Debloating transformations are complex and ensuring they do not compromise the soundness of the program is challenging whether they are implemented pre- (i.e., S2S), during (i.e., CBS), or post-production (i.e., B2B, SL). Flawed or incomplete transformations have the potential to negatively impact programs in many ways that can manifest as logic bugs, run-time errors, crashes, and in some cases, introducing new vulnerabilities. We discovered ample evidence of this during our benchmark performance tests (Section 4.2), where 36.6% of the debloated binaries produced in our evaluation failed to execute their retained functionality.
@@
It is important to note that DIFFER does not and cannot provide formal guarantees of soundness in debloaters or the debloated programs they produce. Like other dynamic analysis testing approaches, DIFFER cannot exhaustively test the input space for complex programs in the general case. Still, DIFFER is quite useful for post-debloating validation as it is user-friendly and only requires a moderate about of user expertise to configure.
Using the general debloating specifications we created in Section 3.3 as a starting point, we configured DIFFER to test both retained and debloated functions for each benchmark program. We ran DIFFER on each of the debloated benchmark programs that successfully completed their performance test (a total of 90 variants) for a maximum of 12 hours to identify crashes, inconsistencies, or errors introduced during debloating as well as failures to remove functionality marked for debloating. Our results are concerning: after filtering for false positives, DIFFER discovered errors or crashes in 27.8% (25 of 90) of debloated benchmarks. Additionally, debloating tools failed to remove features marked for debloating in 60% (39 of 65) of the remaining debloated benchmarks. Of the final 26 debloated benchmarks that passed testing by DIFFER, thirteen were produced by Binary Reduce (static), seven were produced by libfilter, four were produced by LMCAS-SIFT, and OCCAM and CHISEL each produced one. Across our evaluation, 50% (20 of 40) of attempts to remove Type I bloat were ultimately successful, and only 3.3% of attempts to remove Type II bloat were successful (6 of 180) [^4] Five successes were on low-complexity benchmarks, one was on a medium complexity benchmark.. Our results indicate that the general lack of attention paid to post-debloating validation has resulted in over-reporting of successful debloating in the literature. This demonstrates the need for closer attention to post-debloating validation in future research to ensure accurate reporting of debloating success.
@@ 5 Discussion and Key Findings
In this section we discuss our key findings and provide actionable recommendations for future debloating research.
5.1 Low Tool Maturity
Our most important finding is that debloating tools currently lack the maturity required to be used on real-world software. This is evidenced by a slim success rate of 42.5% across all tools and complexity levels at producing debloated programs or libraries that can pass a performance test. The success rate across medium- and high-complexity benchmarks is only 22%, indicating that the current generation of debloaters are suitable for use on lower complexity programs only. To advance the state of the art in debloating, new approaches should focus on supporting more complex software packages and programming language features / paradigms.
5.2 Soundness Issues
When tested for correctness and robustness with DIFFER, only two of the ten tools we evaluated, Binary Reduce static and dynamic, did not produce unsound debloated binaries (i.e., DIFFER did not detect any crashes, failures, or errors in kept functionality). In the case of Binary Reduce (dynamic), this tool produced only one debloated benchmark, and DIFFER did detect that it still contained features marked for debloating. As such, it is possible this binary is only sound because the tool did not successfully debloat any features, which is supported by the fact that this binary increased in size after debloating.
@@ 5.2 Soundness Issues
5.3 Marginal Performance/Security Benefits
Finally, the tools we evaluated demonstrated consistent and significant improvements in two metrics only: binary size and gadget locality. We observed marginal benefits at best for all other performance and security metrics. It is important to note that our security evaluation is limited to code re-usability metrics due to the difficulty of quantifying security improvements. These tools may have had positive impacts such as removing previously-unknown reachable vulnerabilities that are generally not possible to evaluate because they are undecidable.
Considering these marginal benefits in concert with the aforementioned maturity and soundness issues, users are likely to find the benefits of software debloating are currently outweighed by the costs and potential risks in real-world scenarios. Further research is needed to improve the benefits of debloating tools relative to the risks we identified above.